OpenStudy (anonymous):

Which of the following locations has the largest electric field? A, 0.4 cm from a +3.0 μc point charge B. 0.4 cm from a +6.0 μc point charge C. 0.2 cm from a +3.0 μc point charge D. 0.2 cm from a +1.5 μc point charge

OpenStudy (michele_laino):

we have to compute this quantity: \[\Large E\left( r \right) = K\frac{Q}{{{r^2}}}\] for all of your four cases

OpenStudy (anonymous):

ok! how can we do that?

OpenStudy (michele_laino):

for example, first case: \[E\left( r \right) = K\frac{Q}{{{r^2}}} = 9 \times {10^9}\frac{{3 \times {{10}^{ - 6}}}}{{{{\left( {4 \times {{10}^{ - 1}}} \right)}^2}}} = \frac{{270}}{{16}}\]

OpenStudy (anonymous):

and we get 16.875?

OpenStudy (michele_laino):

oops.. I have made an error: \[E\left( r \right) = K\frac{Q}{{{r^2}}} = 9 \times {10^9}\frac{{3 \times {{10}^{ - 6}}}}{{{{\left( {4 \times {{10}^{ - 3}}} \right)}^2}}} = \frac{{27 \times {{10}^9}}}{{16}}\]

OpenStudy (anonymous):

ok! so we get 1687500000?

OpenStudy (michele_laino):

second case: \[E\left( r \right) = K\frac{Q}{{{r^2}}} = 9 \times {10^9}\frac{{6 \times {{10}^{ - 6}}}}{{{{\left( {4 \times {{10}^{ - 3}}} \right)}^2}}} = \frac{{54 \times {{10}^9}}}{{16}}\]

OpenStudy (anonymous):

3375000000?

OpenStudy (michele_laino):

third case: \[E\left( r \right) = K\frac{Q}{{{r^2}}} = 9 \times {10^9}\frac{{3 \times {{10}^{ - 6}}}}{{{{\left( {2 \times {{10}^{ - 3}}} \right)}^2}}} = \frac{{108 \times {{10}^9}}}{{16}}\]

OpenStudy (anonymous):

and then 6750000000 ?

OpenStudy (michele_laino):

fourth case: \[E\left( r \right) = K\frac{Q}{{{r^2}}} = 9 \times {10^9}\frac{{1.5 \times {{10}^{ - 6}}}}{{{{\left( {2 \times {{10}^{ - 3}}} \right)}^2}}} = \frac{{54 \times {{10}^9}}}{{16}}\]

OpenStudy (michele_laino):

what is the highest value?

OpenStudy (anonymous):

the highest value is case 3? with 108*10^9 / 16 ?

OpenStudy (michele_laino):

yes! that's right!

OpenStudy (anonymous):

yAY! what does that mean about the solution? :/ oi am not sure which choice it would be :/

OpenStudy (michele_laino):

it is the third option as you have said before

OpenStudy (anonymous):

ohh okay! i see:) thank you!! onto the next one then :)

OpenStudy (michele_laino):

ok!